HIGH RESOLUTION INFRARED SPECTRA OF ^{13}C-METHANOL IN THE OH-STRETCH REGION

\underline{A. CHIROKOLAVA}, X. WANG, T. J. CRONIN, and D. S. PERRY, Department of Chemistry, University of Akron, Akron, OH 44325-3601; L.-H. XU, Department of Physical Sciences, University of New Brunswick, Saint John, New Brunswick, Canada E2L 4L5.

High-resolution infrared spectra of the ^{13}C-methanol ν_{1} band have been obtained by slit-jet absorption spectroscopy. A total of 1278 lines have been recorded from 3665 to 3708 cm^{−1} with several ”holes” resulting from strong water absorption lines. Thirty five subbands of the ν_{1} band with K = 0, 1, 2 have been assigned based on polynomial fits and ground-state combination differences. Half of the upper states are split by perturbations with matrix elements in the range of 1-3 cm^{−1}. The deperturbed data show an approximately regular pattern of torsional energies with a torsional tunnelling splitting of 7 cm^{−1} at K = 0.